### v2

# Measuring Iterative Temporal Reasoning with [Uncaptioned image]Time Puzzles

Abstract
We introduce [Uncaptioned image]Time Puzzles, [^1] https://github.com/jaaack-wang/Time-Puzzles a constraint-based date inference task for evaluating iterative temporal reasoning. Each puzzle combines factual temporal anchors with (cross-cultural) calendar relations, admits one or multiple valid solution dates, and is algorithmically generated for controlled, dynamic, and continual evaluation. Across 13 diverse LLMs, [Uncaptioned image]Time Puzzles well distinguishes their iterative temporal reasoning capabilities and remains challenging without tools: GPT-5 reaches only 49.3% accuracy and all other models stay below 31%, despite the dataset’s simplicity. Web search consistently yields substantial gains and using code interpreter shows mixed effects, but all models perform much better when constraints are rewritten with explicit dates, revealing a gap in reliable tool use. Overall, [Uncaptioned image]Time Puzzles presents a simple, cost-effective diagnostic for tool-augmented iterative temporal reasoning.
Measuring Iterative Temporal Reasoning with [Uncaptioned image]Time Puzzles
Zhengxiang Wangthanks: Equal contribution. ZW conceived the research idea. ZD designed and implemented the algorithm for dataset generation. ZW executed the LLM experiments. Both authors contributed significantly to drafting and revising the manuscript. Department of Linguistics & IACS Stony Brook University zhengxiang.wang@stonybrook.edu          Zeyu Dong11footnotemark: 1 Department of Applied Math and Statistics Stony Brook University zeyu.dong@stonybrook.edu
1 Introduction
Time is a fundamental attribute of human societies, and temporal reasoning is a core component of human intelligence. In practice, humans rarely reason about time in isolation or by rote memorization; instead, they routinely rely on external tools such as clocks, calendars, written records, and online resources to support temporal inference and verification. With the growing deployment of tool-using, agentic LLMs across consumer-facing, research, and industrial settings (OpenRouter.ai, 2025), it is both timely and important to evaluate iterative temporal reasoning in tool-augmented contexts.
Despite extensive study, existing temporal reasoning benchmarks remain mismatched with how temporal reasoning is performed in practice. Prior work largely evaluates temporal understanding in static, single-shot settings (Ning et al., 2020; Zhou et al., 2021, 2019; Qin et al., 2021). While recent benchmarks extend evaluation to LLMs and situated contexts (Tan et al., 2023; Chu et al., 2024; Wang and Zhao, 2024; Wei et al., 2025; Chen et al., 2021), they still do not explicitly evaluate tool-augmented, iterative temporal reasoning.
To address this gap, we propose [Uncaptioned image]Time Puzzles, a family of simple, puzzle-like constraint-based date inference tasks designed to evaluate precise, iterative temporal reasoning with tools. As illustrated in Figure 1, each puzzle presents natural-language temporal constraints that combine factual temporal anchors with (cross-cultural) calendar-based relations, and asks the model to infer the date(s) that jointly satisfy all constraints. Solving a puzzle requires iteratively proposing, refining, and verifying candidate dates using external tools (e.g., calendars or web search), emphasizing constraint-based reasoning rather than rote memorization or single-shot recall. This formulation reflects real-world settings such as scheduling, planning, and historical analysis, where systems must integrate factual information with calendar structure to produce temporally consistent decisions.
Figure 1: Refer to captionTime Puzzles: a simple date inference task requiring iterative tool-aided temporal reasoning.
Our main contributions are threefold:
  1. 1.
    We introduce [Uncaptioned image]Time Puzzles, a novel and dynamic benchmark for evaluating iterative temporal reasoning with tools. The benchmark is enabled by a controllable generation algorithm that supports continual evaluation beyond static test sets prone to memorization.
  2. 2.
    We conduct a comprehensive evaluation across 13 diverse LLMs. Although tool access improves accuracy, substantial gaps remain in producing fully consistent solutions, highlighting the need for more reliable tool use. We show that [Uncaptioned image]Time Puzzles serves as a practical diagnostic for progress in tool-augmented iterative temporal reasoning.
  3. 3.
    For reproducibility, we release the code and datasets for generating challenging constraint-based time puzzles.
2 [Uncaptioned image]Time Puzzles
This section defines the [Uncaptioned image]Time Puzzles task and describes its algorithmic generation process.
2.1 Task Formulation
Let $\mathcal{D}$ denote the set of all possible Gregorian dates, and let $\mathcal{C}$ map a natural language fact $t$ to the set of dates for which $t$ holds, i.e., $\mathcal{C}(t)\subseteq\mathcal{D}$. As illustrated in Figure 1, each puzzle consists of $N$ natural language facts $F=\{t_{1},t_{2},\dots,t_{N}\}$, with an answer set $\mathcal{A}$ that satisfies all constraints jointly:
$\mathcal{A}=\bigcap_{i=1}^{N}\mathcal{C}(t_{i})$ (1)
We restrict our study to puzzles with non-empty answer sets ($|\mathcal{A}|\geq 1$) to focus on LLMs’ ability to identify valid solution dates via iterative temporal reasoning, and cap $|\mathcal{A}|$ at 6 for exploration.
2.2 Data Generation
Algorithm
We generate [Uncaptioned image]Time Puzzles instances using the dynamic procedure in Algorithm 1. The algorithm takes as input a set of pre-configured temporal fact templates and a target solution size $M$, and returns a puzzle instance $(F,\mathcal{A})$ with $|\mathcal{A}|=M$.
Generation proceeds as a randomized search-and-find process. We first anchor the puzzle with a real-world event description $t_{kb}$ (often requiring web search) and sample a seed date $d_{\text{seed}}$ that is temporally consistent with $t_{kb}$. We then randomly sample 3-5 more temporal constraints (e.g., calendar relations) until a solver yields an answer set of size $M$. To improve data generation efficiency, the solver prunes the search space using an entropy-based approach (see Appendix A.2 for details). For our datasets detailed below, it takes around 2 seconds to generate each puzzle in a single core CPU.
Temporal Facts
To ensure diversity and realism, we create a diverse taxonomy of temporal facts (Table 2 in Appendix A.2) that supports constraints at three granularities: year, month, and day. Our facts span lesser-known historical events, well-known anchors, and calendar relations (e.g., day-of-week and non-Gregorian calendars), among others.
Because [Uncaptioned image]Time Puzzles is designed to probe tool-augmented iterative reasoning, we append to each puzzle a randomly sampled, trivial historical event from a curated list of 50 that potentially require web search. All events occur no later than 2023 and are easily searchable, as manually verified during curation (Appendix A.1). We leave harder-to-find facts for future stress-testing.
Resulting Datasets
We construct two conceptually equivalent datasets for controlled evaluation. The default dataset uses implicit constraints (e.g., an event) to target tool-augmented iterative temporal reasoning. The paired dataset rewrites each implicit constraint by spelling out its relevant dates explicitly, isolating pure iterative temporal reasoning. Each dataset contains 600 puzzles, with solution set sizes evenly distributed from 1 to 6.
Input: Target facts $N$, Target solution size $M$, Universal Date Set $\mathcal{D}$
Output: A puzzle instance $(F,\mathcal{A})$
1
2while True do
 // Step 1: Randomly pick a seed date
3 $t_{kb}\leftarrow\text{SelectRandomKnowledgeBasedEvent()}$;
4 $d_{seed}\leftarrow\text{RandomDate}(\mathcal{C}(t_{kb}))$;
5
 // Step 2: Generate Facts
 // See Appendix A.2 for detail
6 $F\leftarrow\textsc{GenerateFact}(t_{kb},d_{seed},N,\mathcal{D})$;
7
 // Step 3: Solve and Validate
 // See Appendix A.2 for detail
8 $\mathcal{A}\leftarrow\textsc{SolvePuzzle}(F,\mathcal{D})$;
9
 // Step 4: Output condition
10 if $|\mathcal{A}|==M$ then
11    return $(F,\mathcal{A})$;
12
13   end if
14
15 end while
Algorithm 1 TimePuzzles Data Generation Process
Figure 2: Average exact match accuracy across solution counts, with and without web search (only on solution counts 1, 3, 5). GPT-4.1/5 run live web search; for open-weight models we re-use the same cited GPT web results.
3 Experiments
This section first outlines the overall experimental setup and then reports results on [Uncaptioned image]Time Puzzles under both tool-free and tool-augmented settings.
3.1 Experimental Setup
Models
We evaluate 13 diverse LLMs from four model families (Table 1) released in 2025. These include four proprietary GPT models (GPT-5 and GPT-4.1 series) OpenAI (2025b) and nine open-weight models: GPT-oss-20B OpenAI et al. (2025) plus instruction/reasoning variants from DeepSeek-V3.2 DeepSeek-AI et al. (2025), Qwen3 Yang et al. (2025), and Ministral-3-8B Mistral AI (2025). See Table 4 in Appendix C for further details.
Prompting
All models are prompted using zero-shot chain-of-thought (CoT) prompting Kojima et al. (2022) to encourage step-by-step reasoning. The prompt template is provided in Appendix E.
Metrics
Given our emphasis on precise iterative reasoning, we primarily evaluate performance using exact match (EM) between the predicted and gold sets. We also consider F1 score and Jaccard index, which are highly correlated with EM (Spearman’s $r=0.92$) and differ from EM slightly (see Appendix B.2). We additionally report output token counts as a proxy for reasoning efficiency.
3.2 No Tool Use
Table 1 shows the aggregated tool-less results using both the implicit (default) and explicit constraints.
Despite its low construction cost, Time Puzzles is challenging for tool-less LLMs.
Our dataset is a proof of concept built entirely from easy-to-search factual knowledge that predates each model’s training cutoff or release date. Yet even GPT-5 achieves only 49.3% EM on the dataset, with all other models below 31% EM. Providing explicit dates instead of implicit factual constraints improves performance, but a clear reasoning gap remains: GPT-5 reaches the best EM at 80.7% and it takes 685B parameters for DeepSeek V3.2-R to get only 77.0% EM. Moreover, model performance tends to decrease as the number of solutions grow (Figure 2).
Exact Match (%) Output Tokens
Implicit Explicit Implicit Explicit
GPT-5 49.3 80.7 (+31.4) 4382.4 3416.9 (-965.5)
GPT-5-nano 14.0 57.2 (+43.2) 5271.0 4903.9 (-367.1)
GPT-4.1 30.0 58.8 (+28.8) 671.6 593.0 (-78.6)
GPT-4.1-nano 8.2 40.0 (+31.8) 1116.9 1023.6 (-93.3)
GPT-oss-20B 10.2 43.8 (+33.6) 5601.0 4965.2 (-635.8)
DeepSeek-V3.2 27.0 68.5 (+41.5) 785.4 704.6 (-80.8)
DeepSeek-V3.2-R 33.3 77.0 (+43.7) 8276.5 5821.8 (-2454.7)
Qwen3-4B-Inst 2.7 21.2 (+18.5) 2621.0 2563.5 (-57.5)
Qwen3-4B-Think 5.5 38.7 (+33.2) 9529.9 10037.3 (+507.4)
Qwen3-8B 10.0 42.7 (+32.7) 7603.6 7161.4 (-442.2)
Qwen3-14B 11.3 46.5 (+35.2) 4879.9 4782.3 (-97.6)
Ministral-3-8B-Inst 3.7 12.3 (+8.6) 1801.0 1650.5 (-150.5)
Ministral-3-8B-R 3.8 18.7 (+14.9) 2117.6 1749.4 (-368.2)
Table 1: Exact-match accuracy and average output tokens, averaged over six solution counts, for implicit (default) and explicit constraints in the tool-less setting.
Time Puzzles effectively distinguishes nuanced iterative temporal reasoning capabilities.
In terms of EM, larger models consistently outperform their smaller counterparts (e.g., GPT-X vs. GPT-X-nano; Qwen3 variants), and reasoning models uniformly surpass instruction-tuned ones (e.g., DeepSeek/Ministral pairs). However, scale alone is insufficient: despite its smaller size, Qwen3-4B-Think substantially outperforms Ministral-3-8B-R, and GPT-oss-20B similarly surpasses both while using only 3.6B active parameters.
Success on Time Puzzles hinges on sustained, structured temporal reasoning rather than exhaustive generation.
For example, GPT-5 outperforms all other reasoning models under both implicit and explicit constraints while producing shorter outputs (except the low-performing Ministral-3-8B-R). Similarly, larger Qwen3 models perform better with much fewer output tokens.
3.3 Web Search
Setup
We investigate if web search can close the performance gap between the implicit and explicit constraints. We consider six LLMs: GPT-5, GPT-4.1, GPT-oss-20B, DeepSeek-V3.2-R, Qwen3-14B, and Ministral-3-8B-R. We run GPT-5 and GPT-4.1 with live web search to represent two paradigms: static, one-off retrieval (GPT-4.1) and agentic, iterative retrieval (GPT-5). We then re-use the search results used in the responses of these two models as fixed context for the open-weight LLMs. For cost reasons, we only run web search on a subset of our dataset with solution counts 1, 3, and 5.
Results
Figure 2 shows that web search consistently improves performance across different solution counts, but does not close the performance gap between implicit and explicit constraints. Moreover, the marginal gains decrease as the solution count increases. Notably, GPT-5-style web search yields larger improvements than GPT-4.1-style web search, suggesting that iterative temporal reasoning benefits more from agentic, repeated retrieval than from a single retrieval pass.
Overall, these results indicate that even with access to web search, state-of-the-art LLMs still struggle to reliably retrieve and, more importantly, integrate evidence to solve iterative temporal reasoning problems and fully resolve temporal constraints.
Figure 3: Exact-match accuracy (left) and output tokens (right) for GPT-5 (top) and GPT-4.1 (bottom) under different conditions for single-solution puzzles. Numbers annotated on bars are mean values, beside error bars. +Web enables web search. +CI enables Code Interpreter.
3.4 Code Interpreter
Setup
We evaluate whether enabling Code Interpreter (+CI) improves performance and/or reasoning efficiency under the previously discussed conditions. For cost reasons, we test only GPT-5 and GPT-4.1 on single-solution puzzles, representing static and agentic uses of +CI, respectively.
Results
Figure 3 shows that enabling Code Interpreter (+CI) degrades GPT-5 performance on implicit constraints, both with and without web search, whereas GPT-4.1 benefits from +CI, particularly when combined with web search. In contrast, performance on explicit constraints remains largely unchanged for both models. Moreover, +CI consistently reduces GPT-5 output length, but does not yield similar verbosity reductions for GPT-4.1. Overall, +CI is not a reliable substitute for resolving implicit temporal constraints, though it can improve reasoning efficiency and complement web search for capable non-reasoning models.
4 Related Work
Temporal reasoning has long been studied in NLP, including event ordering, implicit event inference, and temporal commonsense reasoning (Ning et al., 2020; Zhou et al., 2021, 2019; Qin et al., 2021). More recent work focuses on diagnosing temporal reasoning in LLMs through broader benchmarks and systematic analyses across diverse task formats (Tan et al., 2023; Chu et al., 2024; Wang and Zhao, 2024; Wei et al., 2025), as well as targeted probes of specific temporal phenomena and synthetic constructions (Islakoglu and Kalo, 2025; Wei et al., 2023; Fatemi et al., 2025; Bhatia et al., 2025).
A complementary line of work studies temporal reasoning in situated and dynamic contexts (Chen et al., 2021; Zhang and Choi, 2021; Kasai et al., 2023). Other studies explore more agentic settings that incorporate memory, external computation, or multimodal inputs (Ge et al., 2025; Saxena et al., 2025). Overall, existing benchmarks largely evaluate temporal reasoning in static or retrieval-oriented settings. In contrast, [Uncaptioned image]Time Puzzles focus on constraint-based date inference that benefits from explicit tool use and requires iterative reasoning, providing a complementary evaluation of tool-augmented, agentic temporal reasoning.
5 Conclusion
We propose [Uncaptioned image]Time Puzzles, a constraint-based date inference task that targets iterative, tool-augmented temporal reasoning. Although puzzles are synthetically generated and easy to verify, they expose persistent failures of current instruction- and reasoning-tuned LLMs to reliably resolve implicit temporal constraints, even with tool access. Overall, [Uncaptioned image]Time Puzzles offers a simple, cost-effective diagnostic for tool-augmented iterative temporal reasoning and can be systematically extended to support more challenging evaluations.
Limitations
Synthetic Dataset Construction
[Uncaptioned image]Time Puzzles is enabled by a controllable generation algorithm, but the resulting constraints are not guaranteed to match real-world temporal reasoning in full complexity. In particular, the naturalness of the constraints can vary, and template-based composition may miss ambiguity, conflicting evidence, or underspecified conditions that arise in practice.
Tool and Model Scope
Our experiments do not exhaust the space of models, tools, or agentic workflows. We evaluate off-the-shelf LLMs and off-the-shelf tools, and we do not attempt to build sophisticated tool-use policies to solve Time Puzzles, given the scope of the study. As a result, our results should be interpreted as a diagnostic of current out-of-the-box capabilities rather than an upper bound.
Solution-Less Puzzles
We restrict evaluation to puzzles with at least one valid solution in order to focus on LLMs’ ability to identify correct dates via iterative temporal reasoning. Although our generation algorithm can produce puzzles with zero solutions, such instances often contain trivial self-contradictions (e.g., incompatible year, month, or day constraints) that require little reasoning and are easily solved by smaller models. Designing non-trivial solution-less puzzles that demand substantive iterative reasoning is left to future work.
Ethical Considerations
Data and privacy.
We do not collect human subjects data. Time Puzzles instances are synthetically generated, and our factual anchors are based on public historical information, with no personally identifiable information.
Appendix A Data Generation
A.1 Factual Anchor Collection
Collection Procedure
We manually curate 50 easy-to-search factual instances based on TEMPREASON Tan et al. (2023), a time-sensitive question answering dataset designed to evaluate temporal reasoning over time–event and event–event relations. Our collection procedure is as follows:
  • We randomly sample 400 questions involving time–event and event–event relations from TEMPREASON. Each instance consists of a factual question (e.g., “Which position did Siim Kallas hold in September 1999?”) and its corresponding answer (e.g., “Minister of Finance”).
  • We prompt gpt-oss-20b to answer the sampled questions and find that it could barely answer these questions.
  • For each retained instance, we reframe the question–answer pair into a declarative statement describing an event, with the temporal information removed. We then use web search to recover the corresponding time point or time range during which the event occurred. When necessary, we further reframe the instance to query a closely related event with an unambiguous temporal answer. We stop once 50 valid instances are collected. All temporal answers are obtained from top-ranked Google search results and manually verified by the authors.
  • We use browser-based ChatGPT-5.2-Auto and Gemini 3 Pro with web search enabled to independently fact-check the complete fact table (provided via screenshots). Minor typographical errors are corrected, and one ambiguous instance is replaced. After revision, both models identify no remaining issues.
Fact-Checking Prompt Instruction
See below.
You are given a table containing 50 historical events. For each event, perform a web search to verify whether the provided time information is accurate.
Column meanings:
- start: The year (or date) when the event first began.
- end: The year (or date) when the event ended.
- multi_year_spans: The range of years during which the event took place.
Task instructions:
- Verify whether the provided time information for each event is correct or incorrect based on reliable sources.
- The time may be given at different levels of precision (e.g., year-only vs. full date).
- Judge correctness only at the precision provided.
- Do not mark a time as incorrect solely because it omits a more specific month or day.
Output requirements (one entry per event):
For each event, report:
- The row index
- The event name
- Whether the provided time is correct or incorrect
- A brief explanation supporting your judgment
Including Well-Known Facts
To balance the corpus, we additionally include 20 widely known factual instances. These cover the presidencies of recent U.S. presidents (since 1945) and the lifespans of well-known public figures (e.g., Kobe Bryant, Albert Einstein).
A.2 Algorithm Implementation Details
As mentioned in Section 2.2, we generate each puzzle by sampling a hidden seed date and selecting $N$ natural-language constraints whose conjunction yields a user-defined number of solutions. To ensure diversity and computational efficiency, our pipeline utilizes a comprehensive taxonomy of temporal facts and sorts constraints by Information Gain (IG) during verification.
Fact Taxonomy and Constraint Levels
To ensure diversity, we define a set of fact categories $\mathcal{K}$. Each category represents a specific type of temporal knowledge (e.g., calendar structure, historical events, astronomical cycles). Furthermore, we assign a Constraint Level $L(t)\in\{\textsc{Year},\textsc{Month},\textsc{Day}\}$ to each fact type, indicating the granularity of the information provided. Table 2 summarizes the fact categories used in [Uncaptioned image]Time Puzzles.
Fact Name Description Level
ExplicitYearFact Explicitly states the year (e.g., “The year is 2025”). Year
DecadeFact Specifies the decade (e.g., “The 1990s”). Year
LeapYearFact States that the year is a leap year. Year
ChineseZodiacFact Specifies the Chinese Zodiac animal. Year
PersonAliveFact States a famous person (e.g., Steve Jobs) was alive. Year
USPresidentFact States a specific US President was in office. Year
EventFact States if it is an Olympic or World Cup year. Year
MonthFact Specifies the month of the year. Month
SeasonFact Specifies the season (e.g., Winter) based on month. Month
LunarMonthFact Specifies the month in the Chinese lunar calendar. Month
WeekdayInMonthFact Specifies the Nth occurrence of a weekday (e.g., “2nd Friday”). Day
DayOfMonthFact Specifies the exact day or if it is the first/last day. Day
WeekdayFact Specifies the day of the week (e.g., “It is Monday”). Day
MultiWeekdayFact Specifies a set of possible weekdays (e.g., ··Mon or Tue”). Day
DayOfMonthRangeFact Specifies if the day is before/after a specific day. Day
KnowledgeBaseEventFact Relates the date to a historical event (same day/month/year), such as the 50 facts collected in Appendix A.1. Various
Table 2: Taxonomy of temporal facts used in generation, categorized by constraint granularity level.
Entropy and Information Gain
To generate efficient solvers and measure the utility of each clue, we define the IG for a fact $t$. Let the entropy of the date space given a set of constraints be the logarithm of the size of the valid date set. The information gain of a fact $t$ is the reduction in entropy it provides:
$IG(t)=\log_{2}(|\mathcal{D}|)-\log_{2}(|\mathcal{C}(t)|)$ (2)
Facts with higher IG narrow the search space more aggressively. For example, an exact date has maximal IG, whereas “It is a Monday” has low IG.
Algorithm Details
We provide the specific procedural logic for generating diverse fact sets and solving the puzzles, as referenced in the main methodology. The fact generation process, detailed in Algorithm 2, ensures diversity by selecting facts from different categories and balanced constraint levels. The solution verification process, detailed in Algorithm 3, optimizes performance by first filtering with facts with high IG.
Input: Anchor $t_{kb}$, Seed $d_{seed}$, Count NN, Universal Date Set $\mathcal{D}$
Output: Set of facts $F$
1 $F\leftarrow\{t_{kb}\}$;
2 $UsedCats\leftarrow\{Cat(t_{kb})\}$;
3
// Iteratively generate N-1 distinct facts
4 for $i\leftarrow 1$ to N1N-1 do
 // Ensure balanced levels
5 $TargetLevel\leftarrow\text{DetermineLevel}(i,N)$;
6 $CandidateCats\leftarrow\{k\in\mathcal{K}\mid k\notin UsedCats,Level(k)==TargetLevel\}$;
7 $k_{new}\leftarrow\text{RandomSelect}(CandidateCats)$;
8 $t_{new}\leftarrow\text{InstantiateFact}(k_{new},d_{seed})$;
9 $F\leftarrow F\cup\{t_{new}\}$;
10 $UsedCats\leftarrow UsedCats\cup\{k_{new}\}$;
11
12 end for
13return $F$;
Algorithm 2 GenerateFact Procedure
Input: Set of facts $F$, Universal Date Set $\mathcal{D}$
Output: Solution set $\mathcal{A}$
// Sort facts to optimize intersection
1 Compute $IG(t)$ for all $t\in F$;
2 $F_{sorted}\leftarrow\text{SortDescending}(F,\text{key}=IG)$;
3
[…]
5 for $t\in F_{sorted}$ do
6 $\mathcal{A}\leftarrow\mathcal{A}\cap\mathcal{C}(t)$;
7
8 end for
9return $\mathcal{A}$;
Algorithm 3 SolvePuzzle Procedure
Appendix B Evaluation Metrics
In the paper, we primarily use exact match accuracy to measure model performance to emphasize our focus on precise iterative reasoning, but we also consider F1 and Jaccard Index to enhance our evaluation. We describe their computation and the empirical difference between exact match accuracy and F1/Jaccard Index.
B.1 Metric Computation
Let $Y_{i}$ denote the gold solution set and $\hat{Y}_{i}$ the predicted set for instance $i$.
Exact Match Accuracy
$\mathrm{EM}_{i}=\mathbb{I}\left[Y_{i}=\hat{Y}_{i}\right]$ (3)
Precision, Recall, and F1
$\mathrm{Precision}_{i}=\frac{|Y_{i}\cap\hat{Y}_{i}|}{|\hat{Y}_{i}|},\quad\mathrm{Recall}_{i}=\frac{|Y_{i}\cap\hat{Y}_{i}|}{|Y_{i}|},$ (4)
$\mathrm{F1}_{i}=\frac{2\,\mathrm{Precision}_{i}\,\mathrm{Recall}_{i}}{\mathrm{Precision}_{i}+\mathrm{Recall}_{i}}.$ (5)
By convention, when $Y_{i}=\hat{Y}_{i}=\emptyset$, all three metrics are set to 1; when exactly one of the sets is empty, all three metrics are set to 0.
Jaccard Index
$\mathrm{Jaccard}_{i}=\begin{cases}1&\text{if }Y_{i}=\hat{Y}_{i}=\emptyset\\ \dfrac{|Y_{i}\cap\hat{Y}_{i}|}{|Y_{i}\cup\hat{Y}_{i}|}&\text{otherwise}\end{cases}$ (6)
Final reported scores are computed via macro-averaging:
$\mathrm{Metric}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{Metric}_{i},$ (7)
where $N$ is the number of evaluation instances.
Mean Std Dev % Equal Spearman
EM Vs F1 -7.07 20.52 87.98 0.92
EM Vs JI -5.40 16.41 87.98 0.92
Table 3: Exact match (EM) accuracy versus F1/Jaccard Index (JI) based on all the experiments ran on this study (over 19,000 data points).
B.2 Exact Match Versus F1/Jaccard Index
Table 3 compares exact match (EM) accuracy with F1 and Jaccard Index (JI). The mean absolute differences between EM and F1 and between EM and JI are 7.07 and 5.40, respectively. The relatively large standard deviations are expected, as EM is binary at the instance level. Nevertheless, in 87.98% of cases, all three metrics agree, indicating that models typically either solve a puzzle completely or fail entirely. Consistent with this observation, Spearman correlations between EM and the other metrics are high ($r=0.92$).
Model Variant used (snapshot / HF repo) Params Release date Knowledge cutoff
GPT-5 OpenAI (2025a) gpt-5-2025-08-07 Not disclosed 2025-08-07 2024-09-30
GPT-5-nano OpenAI (2025a) gpt-5-nano-2025-08-07 Not disclosed 2025-08-07 2024-05-31
GPT-4.1 OpenAI (2025b) gpt-4.1-2025-04-14 Not disclosed 2025-04-14 2024-06-01
GPT-4.1-nano OpenAI (2025b) gpt-4.1-nano-2025-04-14 Not disclosed 2025-04-14 2024-06-01
GPT-oss-20B OpenAI et al. (2025) openai/gpt-oss-20b 20.9B (3.6B active) 2025-08-05 2024-06
DeepSeek-V3.2-R DeepSeek-AI et al. (2025) deepseek-reasoner 685B 2025-12-01 Not disclosed
DeepSeek-V3.2 DeepSeek-AI et al. (2025) deepseek-chat 685B 2025-12-01 Not disclosed
Qwen3-4B-Inst Yang et al. (2025) Qwen/Qwen3-4B-Instruct-2507 4B 2025-08-06 Not disclosed
Qwen3-4B-Think Yang et al. (2025) Qwen/Qwen3-4B-Thinking-2507 4B 2025-08-06 Not disclosed
Qwen3-8B Yang et al. (2025) Qwen/Qwen3-8B 8B 2025-04-29 Not disclosed
Qwen3-14B Yang et al. (2025) Qwen/Qwen3-14B 14B 2025-04-29 Not disclosed
Ministral-3-8B-Inst Mistral AI (2025) mistralai/Ministral-3-8B-Instruct-2512 8B 2025-12-02 Not disclosed
Ministral-3-8B-R Mistral AI (2025) mistralai/Ministral-3-8B-Reasoning-2512 8B 2025-12-02 Not disclosed
Table 4: Model variants and specifications used in our experiments. We identify OpenAI models by their dated snapshot IDs and open-weight models by their Hugging Face repository IDs. Release dates reflect snapshot timestamps or official provider documentation. Note that precise release and knowledge cutoff dates may vary slightly across sources.
Appendix C Models
Inference Details
We evaluate four OpenAI GPT-family models via the OpenAI API (GPT-5, GPT-5-nano, GPT-4.1, and GPT-4.1-nano), using the dated snapshot identifiers shown in Table 4. We also evaluate two DeepSeek-V3.2 endpoints via the DeepSeek API (deepseek-chat and deepseek-reasoner; accessed in December 2025). For open-weight models, we run local inference with vLLM Kwon et al. (2023). We use default generation configuration for all models.
Model Specification
Table 4 lists the exact model variants (API snapshot IDs or Hugging Face repo IDs), parameter sizes, release dates, and (when disclosed) knowledge cutoffs.
Web Search Implementation
We enable web search for GPT-4.1 and GPT-5 via the official Responses API by providing the web search tool at inference time and setting tool choice as required. From each model’s output JSON, we extract the URLs cited in its response and scrape the corresponding webpage contents to construct a fixed web-search context for all open-weight LLMs. This context is prepended to the prompt in the form: ‘‘Here is some web search context that may help you: {scraped contents from cited URLs}.’’
Code Interpreter Implementation
We enable Code Interpreter for GPT-4.1 and GPT-5 via the official Responses API, with tool choice set as required.
Web Search + Code Interpreter Implementation
Providing Web Search and Code Interpreter simultaneously does not guarantee that both tools are invoked during generation, even when tool choice is set as required. To control costs and ensure consistent retrieval, we therefore reuse the web-search context collected from the Web Search–only setting and prepend it to the prompt for GPT models, while enabling Code Interpreter with tool choice set as required.
Appendix D Additional Results
Figure 4 reports average exact match accuracy across solution counts for the remaining models (for which web search is not implemented) under both implicit and explicit constraints.
Figure 4: Average exact-match accuracy across solution counts for other models evaluated without web search.
Appendix E Prompts
We use the following prompt template. The placeholder {constraints} is replaced with the instance-specific constraint set at inference time.
From the time-related constraints below, determine all valid date(s) (if any) that satisfy them. Depending on the conditions, the result may be zero, one, or multiple dates. Unless otherwise stated, interpret all constraints using the Gregorian calendar.
Note: Seasons are defined as:
- Winter: December, January, February
- Spring: March, April, May
- Summer: June, July, August
- Autumn: September, October, November
The constraints are as follows:
{constraints}
Carefully review the constraints and reason step-by-step to identify all valid date(s). After thorough consideration, end your response on a new line with "MY ANSWER: " followed by the valid date(s) in the format "YYYY-$M$-DD". If there are multiple valid dates, list them separated by commas. If no valid date exists, respond with "MY ANSWER: None".

### v3

@@
 Abstract
-We introduce [Uncaptioned image]Time Puzzles, [^1] https://github.com/jaaack-wang/Time-Puzzles a constraint-based date inference task for evaluating iterative temporal reasoning.
-Each puzzle combines factual temporal anchors with (cross-cultural) calendar relations, admits one or multiple valid solution dates, and is algorithmically generated for controlled, dynamic, and continual evaluation.
-Across 13 diverse LLMs, [Uncaptioned image]Time Puzzles well distinguishes their iterative temporal reasoning capabilities and remains challenging without tools: GPT-5 reaches only 49.3% accuracy and all other models stay below 31%, despite the dataset’s simplicity.
-Web search consistently yields substantial gains and using code interpreter shows mixed effects, but all models perform much better when constraints are rewritten with explicit dates, revealing a gap in reliable tool use.
-Overall, [Uncaptioned image]Time Puzzles presents a simple, cost-effective diagnostic for tool-augmented iterative temporal reasoning.
+Tool use, such as web search, has become a standard capability even in freely available large language models (LLMs).
+However, existing benchmarks evaluate temporal reasoning mainly in static, non-tool-using settings, which poorly reflect how LLMs perform temporal reasoning in practice.
+We introduce [Uncaptioned image] Time Puzzles, a constraint-based date inference task for evaluating iterative temporal reasoning with tools.
+Each puzzle combines factual temporal anchors with (cross-cultural) calendar relations and may admit one or multiple valid dates.
+The puzzles are algorithmically generated, enabling controlled and continual evaluation.
+Across 13 LLMs, even the best model (GPT-5) achieves only 55.3% accuracy without tools, despite using easily searchable facts.
+While web search improves performance, models perform substantially better when constraints are rewritten with explicit dates, removing the need for factual lookup.
+These results reveal a gap in reliable tool use for iterative temporal reasoning.
 Measuring Iterative Temporal Reasoning with [Uncaptioned image] Time Puzzles
-Zhengxiang Wangthanks: Equal contribution.
+Zhengxiang Wangthanks: Equal contributions.
 ZW conceived the research idea.
 ZD designed and implemented the algorithm for dataset generation.
-ZW executed the LLM experiments.
+ZW designed and executed the LLM experiments.
 Both authors contributed significantly to drafting and revising the manuscript.
+ZW revised the second version of this draft.
 Department of Linguistics & IACS Stony Brook University zhengxiang.wang@stonybrook.edu          Zeyu Dong11footnotemark: 1 Department of Applied Math and Statistics Stony Brook University zeyu.dong@stonybrook.edu
 1 Introduction
-Time is a fundamental attribute of human societies, and temporal reasoning is a core component of human intelligence.
-In practice, humans rarely reason about time in isolation or by rote memorization; instead, they routinely rely on external tools such as clocks, calendars, written records, and online resources to support temporal inference and verification.
-With the growing deployment of tool-using, agentic LLMs across consumer-facing, research, and industrial settings (OpenRouter.ai, 2025), it is both timely and important to evaluate iterative temporal reasoning in tool-augmented contexts.
-Despite extensive study, existing temporal reasoning benchmarks remain mismatched with how temporal reasoning is performed in practice.
-Prior work largely evaluates temporal understanding in static, single-shot settings (Ning et al., 2020; Zhou et al., 2021, 2019; Qin et al., 2021).
-While recent benchmarks extend evaluation to LLMs and situated contexts (Tan et al., 2023; Chu et al., 2024; Wang and Zhao, 2024; Wei et al., 2025; Chen et al., 2021), they still do not explicitly evaluate tool-augmented, iterative temporal reasoning.
+Modern LLMs are increasingly deployed with external tools, such as web search and code interpreter, even in free customer-facing settings (OpenRouter.ai, 2025).
+These capabilities allow models to retrieve factual information and verify intermediate reasoning steps during inference.
+However, existing temporal reasoning benchmarks largely evaluate models in static, single-shot settings, which do not reflect the iterative, tool-augmented workflows increasingly used in practice.
+Prior work on temporal reasoning has focused on static, single-shot settings (Ning et al., 2020; Zhou et al., 2021, 2019; Qin et al., 2021).
+While recent benchmarks extend evaluation to LLMs and situated contexts (Tan et al., 2023; Chu et al., 2024; Wang and Zhao, 2024; Wei et al., 2025; Chen et al., 2021), they still do not explicitly evaluate tool-augmented, iterative temporal reasoning
 To address this gap, we propose [Uncaptioned image] Time Puzzles, a family of simple, puzzle-like constraint-based date inference tasks designed to evaluate precise, iterative temporal reasoning with tools.
@@
 Figure 1: Refer to caption Time Puzzles: a simple date inference task requiring iterative tool-aided temporal reasoning.
-Our main contributions are threefold:
-  1. 1.
-    We introduce [Uncaptioned image]Time Puzzles, a novel and dynamic benchmark for evaluating iterative temporal reasoning with tools.
-The benchmark is enabled by a controllable generation algorithm that supports continual evaluation beyond static test sets prone to memorization.
-  2. 2.
-    We conduct a comprehensive evaluation across 13 diverse LLMs.
-Although tool access improves accuracy, substantial gaps remain in producing fully consistent solutions, highlighting the need for more reliable tool use.
-We show that [Uncaptioned image]Time Puzzles serves as a practical diagnostic for progress in tool-augmented iterative temporal reasoning.
-  3. 3.
-    For reproducibility, we release the code and datasets for generating challenging constraint-based time puzzles.
+We evaluate 13 diverse LLMs on [Uncaptioned image] Time Puzzles under both tool-less and tool-augmented settings.
+Overall performance remains limited: even the best model (GPT-5) reaches only 55.3% accuracy without tools, despite the use of easily searchable factual anchors.
+While web search consistently improves results, all models perform better when the same puzzles are rewritten to replace factual anchors with explicit dates.
+This gap highlights a key weakness in integrating factual lookup with constraint-based, multi-step temporal reasoning.
 2 [Uncaptioned image] Time Puzzles
@@
 2.1 Task Formulation
-Let $\mathcal{D}$ denote the set of all possible Gregorian dates, and let $\mathcal{C}$ map a natural language fact $t$ to the set of dates for which $t$ holds, i.e., $\mathcal{C}(t)\subseteq\mathcal{D}$.
-As illustrated in Figure 1, each puzzle consists of $N$ natural language facts $F=\{t_{1},t_{2},\dots,t_{N}\}$, with an answer set $\mathcal{A}$ that satisfies all constraints jointly:
+[Uncaptioned image] Time Puzzles is a constraint-based date inference task: given a set of temporal constraints, the goal is to identify the date(s) that jointly satisfy all constraints.
+Formally, let $\mathcal{D}$ denote the set of all possible Gregorian dates, and let $\mathcal{C}$ be an oracle function that maps a natural-language temporal constraint $t$ to the set of dates for which $t$ holds, i.e., $\mathcal{C}(t)\subseteq\mathcal{D}$.
+As illustrated in Figure 1, each puzzle consists of $N$ natural-language constraints $F=\{t_{1},t_{2},\dots,t_{N}\}$.
+The answer set $\mathcal{A}$ contains all dates that satisfy every constraint:
 $\mathcal{A}=\bigcap_{i=1}^{N}\mathcal{C}(t_{i})$ (1)
@@
 2.2 Data Generation
-Algorithm
-We generate [Uncaptioned image]Time Puzzles instances using the dynamic procedure in Algorithm 1.
-The algorithm takes as input a set of pre-configured temporal fact templates and a target solution size $M$, and returns a puzzle instance $(F,\mathcal{A})$ with $|\mathcal{A}|=M$.
-Generation proceeds as a randomized search-and-find process.
-We first anchor the puzzle with a real-world event description $t_{kb}$ (often requiring web search) and sample a seed date $d_{\text{seed}}$ that is temporally consistent with $t_{kb}$.
-We then randomly sample 3-5 more temporal constraints (e.g., calendar relations) until a solver yields an answer set of size $M$.
-To improve data generation efficiency, the solver prunes the search space using an entropy-based approach (see Appendix A.2 for details).
-For our datasets detailed below, it takes around 2 seconds to generate each puzzle in a single core CPU.
-Temporal Facts
-To ensure diversity and realism, we create a diverse taxonomy of temporal facts (Table 2 in Appendix A.2) that supports constraints at three granularities: year, month, and day.
-Our facts span lesser-known historical events, well-known anchors, and calendar relations (e.g., day-of-week and non-Gregorian calendars), among others.
+Input Constraints
+We create a simple yet diverse taxonomy of temporal constraint types (Table 3 in Appendix A.2).
+The taxonomy spans both factual anchors (e.g., historical events, presidents in office, zodiac years) and calendar-structural constraints (e.g., months, seasons, weekdays, or day-of-month relations), each restricting candidate dates at the year, month, or day level.
 Because [Uncaptioned image] Time Puzzles is designed to probe tool-augmented iterative reasoning, we append to each puzzle a randomly sampled, trivial historical event from a curated list of 50 that potentially require web search.
@@
 We leave harder-to-find facts for future stress-testing.
+This paper considers only English as a proof of concept, but [Uncaptioned image] Time Puzzles can be readily adapted to other languages. Figure 1 shows an example puzzle with five constraints.
+Puzzle Generation
+Puzzles are generated algorithmically through randomized constraint sampling.
+Each puzzle combines one of the 50 real-world historical events we manually curated with additional temporal constraints from the taxonomy (e.g., calendar relations).
+We retain combinations that yield puzzles with a desired number of valid solution dates.
+We open-source our code and generated datasets at https://github.com/jaaack-wang/Time-Puzzles.
 Resulting Datasets
@@
 The default dataset uses implicit constraints (e.g., an event) to target tool-augmented iterative temporal reasoning.
-The paired dataset rewrites each implicit constraint by spelling out its relevant dates explicitly, isolating pure iterative temporal reasoning.
+The paired dataset rewrites each implicit constraint by spelling out its relevant dates explicitly, thereby removing the need of factual lookup.
 Each dataset contains 600 puzzles, with solution set sizes evenly distributed from 1 to 6.
-Input: Target facts $N$, Target solution size $M$, Universal Date Set $\mathcal{D}$
-Output: A puzzle instance $(F,\mathcal{A})$
-1
-2while True do
- // Step 1: Randomly pick a seed date
-3 $t_{kb}\leftarrow\text{SelectRandomKnowledgeBasedEvent()}$;
-4 $d_{seed}\leftarrow\text{RandomDate}(\mathcal{C}(t_{kb}))$;
-5
- // Step 2: Generate Facts
- // See Appendix A.2 for detail
-6 $F\leftarrow\textsc{GenerateFact}(t_{kb},d_{seed},N,\mathcal{D})$;
-7
- // Step 3: Solve and Validate
- // See Appendix A.2 for detail
-8 $\mathcal{A}\leftarrow\textsc{SolvePuzzle}(F,\mathcal{D})$;
-9
- // Step 4: Output condition
-10 if $|\mathcal{A}|==M$ then
-11    return $(F,\mathcal{A})$;
-12
-13   end if
-14
-15 end while
-Algorithm 1 TimePuzzles Data Generation Process
+We select 100 puzzles per solution size due to cost constraints, as unrestricted web search is prohibitively expensive, given our limited budget.
+In addition, experiments on a separate set of 600 puzzles yield similar results, supporting the suitability of the current dataset size (see Section 4).
+Quality Control
+Because all constraints are independently verifiable, the resulting puzzles are both scalable and high quality.
+We randomly inspected 30 puzzles, where one author independently attempted to solve them using a web browser and verified the correctness of the generated puzzles.
+This manual inspection also yielded an estimated human baseline accuracy of 80%.
 Figure 2: Average exact match accuracy across solution counts, with and without web search (only on solution counts 1, 3, 5).
@@
 (2025), and Ministral-3-8B Mistral AI (2025).
-See Table 4 in Appendix C for further details.
+See Table 4 in Appendix B for further details.
 Prompting
-All models are prompted using zero-shot chain-of-thought (CoT) prompting Kojima et al. (2022) to encourage step-by-step reasoning. The prompt template is provided in Appendix E.
+All models are prompted using zero-shot chain-of-thought (CoT) prompting Kojima et al. (2022) to encourage step-by-step reasoning. The prompt template is provided in Appendix C.
+Model EM (%) JI (%) F1 (%) # Tks
+GPT-5 55.33 58.62 59.52 4126
+GPT-5-nano 15.83 19.37 20.31 5224
+GPT-4.1 35.17 39.13 40.14 615
+GPT-4.1-nano 9.33 13.53 14.84 1098
+GPT-oss-20B 10.83 13.51 14.32 5704
+DeepSeek-V3.2-R 36.67 39.33 40.08 7862
+DeepSeek-V3.2 30.50 32.89 33.51 737
+Qwen3-4B-Inst 3.00 3.89 4.14 2408
+Qwen3-4B-Think 5.67 6.61 6.89 9701
+Qwen3-8B 10.83 13.45 14.26 7341
+Qwen3-14B 13.17 16.32 17.36 4733
+Ministral-3-8B-Inst 3.50 5.38 6.04 1797
+Ministral-3-8B-R 4.27 4.81 5.04 2058
+Table 1: Exact-match accuracy (EM), Jaccard index (JI), F1 score (F1), and output tokens (Tks), averaged over six solution counts, for the 600 generated puzzles with implicit (default) constraints in the tool-less setting.
 Metrics
-Given our emphasis on precise iterative reasoning, we primarily evaluate performance using exact match (EM) between the predicted and gold sets.
-We also consider F1 score and Jaccard index, which are highly correlated with EM (Spearman’s $r=0.92$) and differ from EM slightly (see Appendix B.2).
-We additionally report output token counts as a proxy for reasoning efficiency.
-3.2 No Tool Use
-Table 1 shows the aggregated tool-less results using both the implicit (default) and explicit constraints.
-Despite its low construction cost, Time Puzzles is challenging for tool-less LLMs.
-Our dataset is a proof of concept built entirely from easy-to-search factual knowledge that predates each model’s training cutoff or release date.
-Yet even GPT-5 achieves only 49.3% EM on the dataset, with all other models below 31% EM.
-Providing explicit dates instead of implicit factual constraints improves performance, but a clear reasoning gap remains: GPT-5 reaches the best EM at 80.7% and it takes 685B parameters for DeepSeek V3.2-R to get only 77.0% EM.
-Moreover, model performance tends to decrease as the number of solutions grow (Figure 2).
-Exact Match (%) Output Tokens
-Implicit Explicit Implicit Explicit
-GPT-5 49.3 80.7 (+31.4) 4382.4 3416.9 (-965.5)
-GPT-5-nano 14.0 57.2 (+43.2) 5271.0 4903.9 (-367.1)
-GPT-4.1 30.0 58.8 (+28.8) 671.6 593.0 (-78.6)
-GPT-4.1-nano 8.2 40.0 (+31.8) 1116.9 1023.6 (-93.3)
-GPT-oss-20B 10.2 43.8 (+33.6) 5601.0 4965.2 (-635.8)
-DeepSeek-V3.2 27.0 68.5 (+41.5) 785.4 704.6 (-80.8)
-DeepSeek-V3.2-R 33.3 77.0 (+43.7) 8276.5 5821.8 (-2454.7)
-Qwen3-4B-Inst 2.7 21.2 (+18.5) 2621.0 2563.5 (-57.5)
-Qwen3-4B-Think 5.5 38.7 (+33.2) 9529.9 10037.3 (+507.4)
-Qwen3-8B 10.0 42.7 (+32.7) 7603.6 7161.4 (-442.2)
-Qwen3-14B 11.3 46.5 (+35.2) 4879.9 4782.3 (-97.6)
-Ministral-3-8B-Inst 3.7 12.3 (+8.6) 1801.0 1650.5 (-150.5)
-Ministral-3-8B-R 3.8 18.7 (+14.9) 2117.6 1749.4 (-368.2)
-Table 1: Exact-match accuracy and average output tokens, averaged over six solution counts, for implicit (default) and explicit constraints in the tool-less setting.
-Time Puzzles effectively distinguishes nuanced iterative temporal reasoning capabilities.
-In terms of EM, larger models consistently outperform their smaller counterparts (e.g., GPT-X vs.
-GPT-X-nano; Qwen3 variants), and reasoning models uniformly surpass instruction-tuned ones (e.g., DeepSeek/Ministral pairs).
-However, scale alone is insufficient: despite its smaller size, Qwen3-4B-Think substantially outperforms Ministral-3-8B-R, and GPT-oss-20B similarly surpasses both while using only 3.6B active parameters.
-Success on Time Puzzles hinges on sustained, structured temporal reasoning rather than exhaustive generation.
-For example, GPT-5 outperforms all other reasoning models under both implicit and explicit constraints while producing shorter outputs (except the low-performing Ministral-3-8B-R).
-Similarly, larger Qwen3 models perform better with much fewer output tokens.
-3.3 Web Search
-Setup
-We investigate if web search can close the performance gap between the implicit and explicit constraints.
-We consider six LLMs: GPT-5, GPT-4.1, GPT-oss-20B, DeepSeek-V3.2-R, Qwen3-14B, and Ministral-3-8B-R.
-We run GPT-5 and GPT-4.1 with live web search to represent two paradigms: static, one-off retrieval (GPT-4.1) and agentic, iterative retrieval (GPT-5).
-We then re-use the search results used in the responses of these two models as fixed context for the open-weight LLMs.
-For cost reasons, we only run web search on a subset of our dataset with solution counts 1, 3, and 5.
-Results
-Figure 2 shows that web search consistently improves performance across different solution counts, but does not close the performance gap between implicit and explicit constraints.
-Moreover, the marginal gains decrease as the solution count increases.
-Notably, GPT-5-style web search yields larger improvements than GPT-4.1-style web search, suggesting that iterative temporal reasoning benefits more from agentic, repeated retrieval than from a single retrieval pass.
-Overall, these results indicate that even with access to web search, state-of-the-art LLMs still struggle to reliably retrieve and, more importantly, integrate evidence to solve iterative temporal reasoning problems and fully resolve temporal constraints.
-Figure 3: Exact-match accuracy (left) and output tokens (right) for GPT-5 (top) and GPT-4.1 (bottom) under different conditions for single-solution puzzles.
+Let $Y_{i}$ denote the gold solution set and $\hat{Y}_{i}$ the predicted set for instance $i$. We compute exact match (EM), F1 (F1), and Jaccard index (JI):
+$EM_{i}=\mathbb{I}[Y_{i}=\hat{Y}_{i}].$ (2)
+$P_{i}=\tfrac{|Y_{i}\cap\hat{Y}_{i}|}{|\hat{Y}_{i}|},\;R_{i}=\tfrac{|Y_{i}\cap\hat{Y}_{i}|}{|Y_{i}|},\;F1_{i}=\tfrac{2P_{i}R_{i}}{P_{i}+R_{i}}.$ (3)
+$JI_{i}=\begin{cases}1&Y_{i}=\hat{Y}_{i}=\emptyset\\ \tfrac{|Y_{i}\cap\hat{Y}_{i}|}{|Y_{i}\cup\hat{Y}_{i}|}&\text{otherwise}.\end{cases}$ (4)
+Given our emphasis on precise iterative reasoning, we treat EM as the primary metric.
+Tools
+We consider the following tool-augmented settings: web search, Code Interpreter (+CI), and their combinations.
+Due to cost constraints, these experiments are conducted on puzzles with solution counts 1, 3, and 5 and with selected models.
+More specifically, for web search, we test six best-performing LLMs across different model families (GPT-5, GPT-4.1, GPT-oss-20B, DeepSeek-V3.2-R, Qwen3-14B, and Ministral-3-8B-R), whereas for +CI, we only evaluate GPT-5 and GPT-4.1 both with and without web search.
+We use off-the-shelf tools via OpenAI Response API and we include GPT-5 and GPT-4.1 across all tool-augmented settings to represent tool-using capablities by a reasoning model (GPT-5) as well as an instruction model (GPT-4.1).
+Note that GPT-5 and GPT-4.1 are run with live web search and their retrieved results are reused as fixed context for the open-weight models.
+3.2 Results
+Despite its simplicity, [Uncaptioned image] Time Puzzles is challenging for tool-less LLMs.
+Our dataset is a proof of concept built from easy-to-search facts that predates each model’s training cutoff or release date.
+Yet even GPT-5 achieves only 55.3% EM on the dataset, with all other models below 37% EM (Table 1).
+JI and F1 yield slightly higher scores, but the overall performance patterns remain consistent, with a Spearman correlation of $r=0.93$.
+Moreover, the three metrics agree in 88.6% of cases, indicating that LLMs typically either solve a puzzle completely or fail to solve it at all.
+Tool use, particularly web search, can greatly improve model performance, but a clear reasoning gap remain.
+Figure 2 shows that web search consistently improves performance across different solution counts, but models nearly always perform the best when the implicit constraints in [Uncaptioned image] Time Puzzles are rewritten with explicit dates to remove the need for factual lookup.
+Similarly, Figure 3 shows that enabling Code Interpreter (+CI) does not close the gap either: +CI degrades GPT-5 performance on implicit constraints, both with and without web search.
+While GPT-4.1 benefits from +CI when combined with web search, the performance is below that of explicit constraints.
+Figure 3: Average EM accuracy (%) for GPT-5 (left) and GPT-4.1 (right) under different conditions across three solution counts (1, 3, and 5).
 Numbers annotated on bars are mean values, beside error bars.
@@
 +CI enables Code Interpreter.
-3.4 Code Interpreter
-Setup
-We evaluate whether enabling Code Interpreter (+CI) improves performance and/or reasoning efficiency under the previously discussed conditions.
-For cost reasons, we test only GPT-5 and GPT-4.1 on single-solution puzzles, representing static and agentic uses of +CI, respectively.
-Results
-Figure 3 shows that enabling Code Interpreter (+CI) degrades GPT-5 performance on implicit constraints, both with and without web search, whereas GPT-4.1 benefits from +CI, particularly when combined with web search.
-In contrast, performance on explicit constraints remains largely unchanged for both models.
-Moreover, +CI consistently reduces GPT-5 output length, but does not yield similar verbosity reductions for GPT-4.1.
-Overall, +CI is not a reliable substitute for resolving implicit temporal constraints, though it can improve reasoning efficiency and complement web search for capable non-reasoning models.
+4 Analysis
+Section 3 shows that [Uncaptioned image] Time Puzzles provides a challenging and cost-effective benchmark for iterative temporal reasoning.
+This section further demonstrates that as a benchmark, [Uncaptioned image] Time Puzzles is both discriminative and reliable.
+[Uncaptioned image] Time Puzzles effectively distinguishes nuanced iterative temporal reasoning capabilities.
+Table 1 clearly shows that larger models consistently outperform their smaller counterparts (e.g., GPT-X vs.
+GPT-X-nano; Qwen3 variants), and reasoning models uniformly surpass instruction-tuned ones (e.g., DeepSeek/Ministral pairs).
+However, scale alone is insufficient: despite its smaller size, Qwen3-4B-Think substantially outperforms Ministral-3-8B-R, and GPT-oss-20B similarly surpasses both while using only 3.6B active parameters.
+Moreover, for reasoning models, higher performance typically uses less output tokens, suggesting that success on [Uncaptioned image] Time Puzzles hinges on sustained, structured reasoning rather than exhaustive generation (e.g., Qwen3 models in Table 1).
+Our findings are robust to prompt variations and puzzle sets.
+Table 2 shows that inserting a detailed step-by-step reasoning instruction into the default prompt (see Appendix C) decreases performance, particularly for larger models.
+Furthermore, repeating the experiments from Section 3 on 600 newly generated puzzles yields negligibly different exact-match (EM) scores, validating the suitability of our evaluation dataset size.
+Type Setting G4.1-Nano G5-Nano G4.1 G5
+Implicit Default 9.3 15.8 35.2 55.3
++Reason Inst 8.8 14.2 25.3 47.3
+New Runs 11.0 16.8 29.8 50.1
+Explicit Default 41.8 60.2 63.0 84.7
++Reason Inst 35.5 55.5 53.5 77.7
+New Runs 42.7 60.2 63.7 80.5
+Table 2: EM accuracy (%) averaged over six solution counts under different settings for both explicit and implicit constraint across the four GPT models.
+Default setting is based on Section 3.
+“+Reason Inst” enhances the default setting with detailed step-by-step reasoning instruction.
+“New Runs” reruns the default setting with completely new puzzles generated from scratch.
 5 Related Work
@@
 We propose [Uncaptioned image] Time Puzzles, a constraint-based date inference task that targets iterative, tool-augmented temporal reasoning.
-Although puzzles are synthetically generated and easy to verify, they expose persistent failures of current instruction- and reasoning-tuned LLMs to reliably resolve implicit temporal constraints, even with tool access.
-Overall, [Uncaptioned image]Time Puzzles offers a simple, cost-effective diagnostic for tool-augmented iterative temporal reasoning and can be systematically extended to support more challenging evaluations.
+Although puzzles are synthetically generated and easy to verify, they are challenging for tool-less LLMs.
+They also expose persistent failures of current instruction- and reasoning-tuned LLMs to reliably resolve implicit temporal constraints, even with tool access.
+Overall, [Uncaptioned image] Time Puzzles offers a simple, cost-effective, and discriminative diagnostic for tool-augmented iterative temporal reasoning and can be systematically extended to support more challenging evaluations or additional languages (e.g., via translation).
 Limitations
@@
 Designing non-trivial solution-less puzzles that demand substantive iterative reasoning is left to future work.
+Language Diversity
+Our work currently only considers English language, due to budget constaints.
+Future studies can extend our study by adapting our evaluation framework to other languages.
+This can be easily done by, for example, converting the puzzle template into other languages.
 Ethical Considerations
@@
 A.2 Algorithm Implementation Details
-As mentioned in Section 2.2, we generate each puzzle by sampling a hidden seed date and selecting $N$ natural-language constraints whose conjunction yields a user-defined number of solutions.
+We generate each puzzle by sampling a hidden seed date and selecting $N$ natural-language constraints whose conjunction yields a user-defined number of solutions.
 To ensure diversity and computational efficiency, our pipeline utilizes a comprehensive taxonomy of temporal facts and sorts constraints by Information Gain (IG) during verification.
@@
 Algorithm 2 SolvePuzzle Procedure
-Appendix B Evaluation Metrics
-In the paper, we primarily use exact match accuracy to measure model performance to emphasize our focus on precise iterative reasoning, but we also consider F1 and Jaccard Index to enhance our evaluation.
-We describe their computation and the empirical difference between exact match accuracy and F1/Jaccard Index.
-B.1 Metric Computation
-Let $Y_{i}$ denote the gold solution set and $\hat{Y}_{i}$ the predicted set for instance $i$.
-Exact Match Accuracy
-$\mathrm{EM}_{i}=\mathbb{I}\left[Y_{i}=\hat{Y}_{i}\right]$ (3)
-Precision, Recall, and F1
-$\mathrm{Precision}_{i}=\frac{|Y_{i}\cap\hat{Y}_{i}|}{|\hat{Y}_{i}|},\quad\mathrm{Recall}_{i}=\frac{|Y_{i}\cap\hat{Y}_{i}|}{|Y_{i}|},$ (4)
-$\mathrm{F1}_{i}=\frac{2\,\mathrm{Precision}_{i}\,\mathrm{Recall}_{i}}{\mathrm{Precision}_{i}+\mathrm{Recall}_{i}}.$ (5)
-By convention, when $Y_{i}=\hat{Y}_{i}=\emptyset$, all three metrics are set to 1; when exactly one of the sets is empty, all three metrics are set to 0.
-Jaccard Index
-$\mathrm{Jaccard}_{i}=\begin{cases}1&\text{if }Y_{i}=\hat{Y}_{i}=\emptyset\\ \dfrac{|Y_{i}\cap\hat{Y}_{i}|}{|Y_{i}\cup\hat{Y}_{i}|}&\text{otherwise}\end{cases}$ (6)
-Final reported scores are computed via macro-averaging:
-$\mathrm{Metric}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{Metric}_{i},$ (7)
-where $N$ is the number of evaluation instances.
-Mean Std Dev % Equal Spearman
-EM Vs F1 -7.07 20.52 87.98 0.92
-EM Vs JI -5.40 16.41 87.98 0.92
-Table 3: Exact match (EM) accuracy versus F1/Jaccard Index (JI) based on all the experiments ran on this study (over 19,000 data points).
-B.2 Exact Match Versus F1/Jaccard Index
-Table 3 compares exact match (EM) accuracy with F1 and Jaccard Index (JI).
-The mean absolute differences between EM and F1 and between EM and JI are 7.07 and 5.40, respectively.
-The relatively large standard deviations are expected, as EM is binary at the instance level.
-Nevertheless, in 87.98% of cases, all three metrics agree, indicating that models typically either solve a puzzle completely or fail entirely.
-Consistent with this observation, Spearman correlations between EM and the other metrics are high ($r=0.92$).
 Model Variant used (snapshot / HF repo) Params Release date Knowledge cutoff
@@
 Note that precise release and knowledge cutoff dates may vary slightly across sources.
-Appendix C Models
+Appendix B Models
 Inference Details
@@
 Providing Web Search and Code Interpreter simultaneously does not guarantee that both tools are invoked during generation, even when tool choice is set as required.
-To control costs and ensure consistent retrieval, we therefore reuse the web-search context collected from the Web Search–only setting and prepend it to the prompt for GPT models, while enabling Code Interpreter with tool choice set as required.
-Appendix D Additional Results
-Figure 4 reports average exact match accuracy across solution counts for the remaining models (for which web search is not implemented) under both implicit and explicit constraints.
-Figure 4: Average exact-match accuracy across solution counts for other models evaluated without web search.
-Appendix E Prompts
-We use the following prompt template. The placeholder {constraints} is replaced with the instance-specific constraint set at inference time.
+To control costs and ensure consistent retrieval, we therefore reuse the web-search context collected from the Web Search-only setting and prepend it to the prompt for GPT models, while enabling Code Interpreter with tool choice set as required.
+Appendix C Prompts
+C.1 Default Prompt
+We use the following prompt template in the main experiments in Section 3. The placeholder {constraints} is replaced with the instance-specific constraint set at inference time.
 From the time-related constraints below, determine all valid date(s) (if any) that satisfy them.
@@
 If no valid date exists, respond with "MY ANSWER: None".
+C.2 Prompt with Step-by-step Reasoning Instruction
+The following prompt is used in the follow-up experiment in Section 4, which builds on the default prompt by inserting detailed step-by-step reasoning instruction.
+From the time-related constraints below, determine all valid date(s) (if any) that satisfy them.
+Depending on the conditions, the result may be zero, one, or multiple dates.
+Unless otherwise stated, interpret all constraints using the Gregorian calendar.
+Carefully follow these step-by-step reasoning instructions:
+1) Read and restate the goal
+- You must find ALL Gregorian calendar date(s) that satisfy EVERY constraint.
+- The answer can be none, one, or many dates.
+2) Normalize definitions you must use
+- Seasons:
+* Winter = Dec, Jan, Feb
+* Spring = Mar, Apr, May
+* Summer = Jun, Jul, Aug
+* Autumn = Sep, Oct, Nov
+- If constraints mention weekday, leap year, end of month, etc., use standard Gregorian rules.
+3) Extract constraints into a checklist
+- Parse the text and list each constraint as a separate item.
+- If any constraint contains an OR, split into separate branches.
+- If constraints contain ranges (e.g., "between 2010 and 2015"), record inclusivity/exclusivity exactly as stated.
+4) Determine the search space (candidate window)
+- Identify the tightest restriction on:
+* Year(s) (specific year, range, century, decade)
+* Month(s) (specific month(s), season, quarter)
+* Day(s) (day number, end of month, first Monday, etc.)
+- Use these to form an initial set of candidate dates or a candidate-generating rule.
+5) Convert each constraint into a filter function
+For each constraint, translate it into a precise test such as:
+- Year filter (e.g., year = 2024; year in [1990, 1999])
+- Month/season filter (e.g., month in {{6,7,8}})
+- Day-of-month filter (e.g., day = 1; day >= 28)
+- Weekday filter (e.g., weekday = Monday)
+- Relative-date relations (e.g., "the day after X", "two weeks before Y")
+- Calendar properties:
+* Leap year rule: leap if divisible by 4, except centuries not divisible by 400.
+* Month lengths: 30/31 days; February 28 or 29 depending on leap year.
+6) Generate candidates in the most restrictive way
+- Prefer generating candidates from the most specific constraints first:
+* If a specific date is implied, compute it directly.
+* If a specific year + season is given, iterate only those months.
+* If a weekday pattern exists ("first Monday of May"), generate those directly instead of brute-forcing all days.
+7) Apply constraints systematically
+- For each candidate date, check constraints one by one.
+- Eliminate candidates immediately when one constraint fails.
+- Keep a running list of remaining candidates.
+8) Handle OR and branching logic correctly
+- If constraints are like: (A OR B) AND C:
+* Compute candidates satisfying A and C
+* Compute candidates satisfying B and C
+* Take the union of those results (deduplicate).
+9) Check for hidden contradictions early
+- Before enumerating too much, look for impossibilities, e.g.:
+* Month is February but day is 30
+* Season says Winter but month says July
+* Date required to be both Monday and Tuesday
+* Before 2000 AND after 2010
+- If a contradiction is found, conclude no solution.
+10) Ensure completeness (dont miss valid dates)
+- Confirm you explored all years/months allowed by the constraints.
+- If you used branching, ensure all branches were evaluated.
+11) Sort and format final answers
+- Sort valid dates chronologically.
+- Output dates in YYYY-$M$-$D$ format.
+- If multiple, separate by commas and a space.
+The constraints are as follows:
+{constraints}
+After thorough consideration, end your response on a new line with "MY ANSWER: "
+followed by the valid date(s) in the format "YYYY-$M$-DD".
+If there are multiple valid dates, list them separated by commas.
+If no valid date exists, respond with "MY ANSWER: None".